1) In an electroplating cell, 0.485g of a metal M are plated out from an acidic solution of MO3 in exactly one hour using a current of 1.50A. what is the molar mass of M?

2) A water electrolysis cell connected in series to the above cell liberated 754cm3 of hydrogen when the temperature was 288K and the atmospheric pressure was 763.8mmHg. The hydrogen was collected over an aqueous solution as in this experiment. The difference in liquid levels inside and outside the eudiometer tube was measured to be 152mm. use this data to calculate the molar mass of M.
1
Expert's answer
2020-11-12T06:31:00-0500

1) "Q= it"

Q=1.5×60×60= 5400c

If 0.485g is electroplated by 5400c, then 96500c can electroplated;

(96500/5400)×0.485

Molar mass=8.67g

2)V= 754cm3

P=763.8 mmhg

T= 288k

"Pv = nRT"

763.8×0.000754 = n× 8.314×288

n=0.00024moles